ADNAN MUHAMMED

Grade 12,

IF A+B+C=180,FIND THE MINIMUM VALUE OF SIN A+SIN B+SIN C?

IF A+B+C=180,FIND THE MINIMUM VALUE OF SIN A+SIN B+SIN C?
 

Grade:11

1 Answers

Abdul Samad
6 years ago
Let A=0 B=0 C=180 Hence SIN A+SIN B+SIN C = 0+0+0 hence minimum value for SIN A+ SIN B+SIN C = 0+0+0=0 According to me 0 is the answer.